# Graph Neural Networks are Heuristics

### v2

@@ 1 Introduction
Heuristics are the grammar of problem-solving: the means by which intelligence—human or artificial—moves from paralysis in the face of combinatorial explosion to progress through approximation. At their core, heuristics are not mere tricks or shortcuts, but strategies for reasoning under constraint, where exact solutions are unattainable or impractical.
The Travelling Salesman Problem (TSP) has long served as a proving ground for such strategies. Exact solvers like Concorde (Applegate et al., 2003) achieve optimality at exponential cost, while carefully engineered heuristics such as Lin-Kernighan-Helsgaun (Lin and Kernighan, 1973; Helsgaun, 2000) embody layers of local improvement rules that deliver state-of-the-art approximations. Simpler constructive approaches (nearest neighbor, farthest insertion) reveal the essential tradeoff between efficiency and accuracy that has always defined heuristic reasoning.
@@
We take a different view: heuristics need not be hand-crafted rules or search procedures. They can emerge directly from structure and learning. Building on the Structure As Search framework (Min et al., 2023; Min and Gomes, 2023, 2025), we formulate TSP as direct permutation learning. Our non-autoregressive (NAR), generative adversarial method produces Hamiltonian cycles without supervision, explicit search, or sequential decoding. The key insight is that TSP’s inherent structure—shortest Hamiltonian cycles—constrains the solution space sufficiently for neural networks to learn effective solutions through structure alone. We employ Gumbel-Sinkhorn relaxation during training and Hungarian algorithm decoding at inference, enabling end-to-end optimization directly from the combinatorial objective.
In this sense, our method constitutes a new form of heuristic. It is not designed by hand, but arises from adversarial training dynamics and structural inductive bias. Like classical heuristics, it does not guarantee optimality; yet, like the best of them, it produces consistently strong solutions across instance sizes. Our results show that heuristics can now be learned rather than engineered, pointing toward a paradigm where structural constraints and generative models replace explicit search as the foundation of combinatorial optimization.

### v3

@@ 1 Introduction
Heuristics are the grammar of problem-solving: the means by which intelligence—human or artificial—moves from paralysis in the face of combinatorial explosion to progress through approximation. At their core, heuristics are not mere tricks or shortcuts, but strategies for reasoning under constraint, where exact solutions are unattainable or impractical. The Travelling Salesman Problem (TSP) has long served as a proving ground for such strategies. Exact solvers like Concorde (Applegate et al., 2003) achieve optimality at exponential cost, while carefully engineered heuristics such as Lin-Kernighan-Helsgaun (Lin and Kernighan, 1973; Helsgaun, 2000) embody layers of local improvement rules that deliver state-of-the-art approximations. Simpler constructive approaches (nearest neighbor, farthest insertion) reveal the essential tradeoff between efficiency and accuracy that has always defined heuristic reasoning.
@@
We take a different view: heuristics need not be hand-crafted rules or search procedures. They can emerge directly from structure and learning. Building on the Structure As Search framework (Min et al., 2023; Min and Gomes, 2023, 2025), we formulate TSP as direct permutation learning. Our non-autoregressive (NAR) method generates Hamiltonian cycles without supervision, explicit search, or sequential decoding. The key insight is that TSP’s inherent structure—shortest Hamiltonian cycles—constrains the solution space sufficiently for neural networks to learn effective solutions. We employ Gumbel-Sinkhorn relaxation during training and Hungarian algorithm decoding at inference, enabling end-to-end optimization. In this sense, our method constitutes a new form of heuristic. It is not designed by hand, but arises from learning from instances. Like classical heuristics, it does not guarantee optimality; yet, like the best of them, it produces consistently strong solutions across instance sizes. Our results show that heuristics can now be learned rather than engineered, pointing toward a paradigm where structural constraints and generative models replace explicit search as the foundation of combinatorial optimization, our implementation can be found at GNNs-are-Strong-Heuristics.